OpenStudy (anonymous):

PLEASE HELP!!! The parking lot at George's Grocery is 60 ft long and 40 ft wide. He plans to expand the width and length the same amount (x). Write an expression in terms of x to represent the new length. Write a second expression in terms of x to represent the new width.

OpenStudy (p0sitr0n):

new length = length + x = 60 +x new width = width + x = 40 + x ?

OpenStudy (anonymous):

Thank you! @P0sitr0n